OpenStudy (anonymous):

convert to rectangular form r=4sec theta

OpenStudy (anonymous):

\[ r = 4 \sec (\theta)= \frac 4 {\cos(\theta)} \\ r \cos(\theta) = 4\\ x =4 \] It is a line

OpenStudy (anonymous):

Would you mind if I asked you more of these problems?

OpenStudy (anonymous):

Please do

OpenStudy (anonymous):

r^2 = 8 sec theta csc theta

OpenStudy (anonymous):

r= 7 csc theta r=12/ 8 csc theta + 11 sin theta theta = - pie/3 theta = pie / 4

OpenStudy (anonymous):

r= 8 sin theta

OpenStudy (anonymous):

r = 7 csc(theta) rsin(θ)=7 y=7

OpenStudy (anonymous):

r=12/ 8 csc theta + 11 sin theta \[r = \frac {12} 8 \frac 1{\sin(\theta)}+ 11 \sin(\theta)\\ r\sin(\theta) = \frac {12} 8 + 11 \sin^2(\theta )\\ y = \frac {12} 8 +11 \frac {x^2}{ r^2}\\ y = \frac {12} 8 +11 \frac {x^2}{ x^2+ y^2}\\ \]

OpenStudy (anonymous):

You should be able to mimic what I did to do the rest. Try them and let me know.

OpenStudy (anonymous):

r= 8 sin theta did you get y= -8

OpenStudy (anonymous):

theta = - pie/3 theta = pie / 4 I'm a bit lost on these two

OpenStudy (anonymous):

Hey, do u by any chance know how to give 3 different polar points that are equivalent to: (0,-5)

OpenStudy (anonymous):

r= 8 sin theta this is not y=8 try to multiply both sides by r thenn \[ r^2= 8 r \sin(\theta) \] In ploar coordinates \[ x= r \cos(\theta) \\ y= r \sin(\theta)\\ x^2+y^2 = r^2 \] Use the above to finish the problem.

OpenStudy (anonymous):

did you get zero for both

OpenStudy (anonymous):

You should get \[ x^2 + y^2 = 8 \sin(\theta)\]

OpenStudy (anonymous):

What does it mean that the angle is constant,like \[\theta = \frac \pi 4 \] It means that we are on a line that makes an angle of 45 degrees with the x-axis. What is the equation of this line?

OpenStudy (anonymous):

Do some practice on http://moltest.missouri.edu/mucgi-bin/calculus.cgi and choose Calculus(II) Parametric and Polar coordinates. You will be good at doing this kind of problems and more.